3n-2=5n+12
3n-5n=12+2
-2n=14
n=-7
check:
3(-7)- 2 = 5(-7) + 12
-23= -23

Answer by annika21(5)

3n-2=5n+12

3n-5n=12+2
-2n=14
n=-7
check:
3(-7)- 2 = 5(-7) + 12
-23= -23